5.) A 1000kg The coefficient of friction is .2 How far does the car travel uphill before it comes to rest? car travelling at 20m/s encounters a 200 incline. Distance- booka
Answer #1

from the free body diagram we have

in y direction we have

Fy=may

Fn-mgcosTheta=0

so Fn=mgcosTheta...........1)

and Fr=muFn

so using eqn 1 we have

Fr=mumgcosTheta........2)

we have the formula for kinetic energy

K.E=1/2mv2

and the work done by the friction force is

Fr*d=mumgcosTheta*d

when the block reaches the top, the kinetic energy changes to potential energy

so P.E=mgh=mgsinTheta*d

P.E=K.E-work done by friction

mgsinTheta*d=1/2mv2-mumgcosTheta*d

gsinTheta*d=1/2v2-mugcosTheta*d

1/2v2=d(gsinTheta+mugcosTheta)

1/2*202=d(9.8*sin20o+0.2*9.8*cos20o)

200=d(5.1936)

d=200/5.1936=38.5m

so answer is 38.5m

PrD 74 f n